# Probing Light Bosonic Dark Matter with Transmon Qubits

### v2

@@ Abstract
In this paper, we investigate the constraints of the transmon qubit, an improved version of the charge qubit, on bosonic light dark matters. Phonon excitations induced by the scattering or absorption of dark matter on a superconductor may destroy the Cooper pair, leading to the production of quasiparticles made by the electron. By measuring the production rate of the quasiparticle density, one may read out the coupling between dark matter and ordinary matter, assuming that these quasiparticles are solely induced by dark matter interactions. For the first time, we show constraints on the parameter space of the dark photon, light scalar dark matter, and axion-like particles from the measurement of quasiparticles in transmon qubit experiments. This study offers insights for the development of quantum qubit experiments aimed at the direct detection of dark matter in underground laboratories.
@@ I Introduction
In recent years, direct detection technologies for DM Billard et al. (2022); Schumann (2019) have developed rapidly. The exclusion limits provided by experiments utilizing nuclear recoil technology have been continuously strengthened, and numerous experiments employing electronic excitation have been conducted to probe DM, enabling the detection of lighter DM candidates. However, these experiments are mostly limited to non-relativistic DM candidates with a mass above the MeV scale. Furthermore, when considering the DM absorption process, the detection threshold can be lowered to the electroVolt (eV) level. In fact, there are also sub-eV DM models with strong theoretical motivations, such as the axion and the dark photon. These particles can be produced in the early universe through mechanisms like the misalignment mechanism and parametric resonance, etc. Direct Detection of ultralight DM Hui et al. (2017) is an important unresolved issue, and various proposals have been put forward to explore sub-eV DM. Electronic excitations in targets with small excitation gaps can be applied to the detection of light DM Essig et al. (2017). The expected target materials include Dirac materials Hochberg et al. (2018), superconductors Hochberg et al. (2016), and narrow-gap semiconductors Hochberg et al. (2017). The emergence of collective excitations like phonon Kahn and Lin (2022) and magnon Trickle et al. (2020), resulting from DM-matter scattering or DM absorption within condensed matter material, represents a promising avenue for DM direct detection. This is particularly relevant given that these excitations possess energies within the ${\cal O}(1\sim 100)$ meV range, aligning well with the energy scales characteristic of light DM. In brief, researchers are actively exploring and employing a variety of methods to detect DM, yet no definitive detection has been achieved. Utilizing new materials or technologies to detect light DM presents an increasingly recognized challenge in the fields of high-energy physics and precision measurement physics. The path to the direct detection of DM remains a lengthy and challenging journey.
@@
In this study, we delve into the potential constraints imposed on light dark matter by superconducting qubits Dixit et al. (2021). These quantum qubits, constructed from superconducting materials and maintained at ultra-low temperatures, have garnered significant interest among quantum computing scientists for their enhanced coherence and stability. Recently, transmon qubits Ristè et al. (2013) as DM detectors have been proposed. Refs. Chen et al. (2023, 2024) have showed how to detect hidden photon and the QCD axion using the direct excitation of transmon qubits. Refs. Das et al. (2024a, b) have shown constraint on the DM-nucleon scattering cross sections using the latest transmon qubit measurements Dixit et al. (2021). In addition, a transmon qubit prototype for DM detection experiment has been designed in Ref. Moretti et al. (2024). We focus on the case of the absorbable DM. Light bosonic DM can be absorbed by the superconducting material into phonons via DM-proton or DM-neutron interactions and the absorption rate can be calculated utilizing the effective field theory developed in Ref. Mitridate et al. (2024). These phonon can break Cooper pairs whenever their energy is above meV, leading to to the production of quasiparticles. With the help of transmon qubit measurements Dixit et al. (2021), we put constraints on the DM couplings with results shown in Figs. 1 and 2, where constrains on the dark photon, light scalar DM and axion-like particle DM are presented. These results are the first solid constraints on light bosonic DM from quantum qubits experiments, and will insights for the development of quantum qubit experiments aimed at the direct detection of dark matter in underground laboratories.
@@ II Light Bosonic DMs
${\cal L}_{\rm DP}\supset-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}X_{\mu\nu}X^{\mu\nu}+\frac{\varepsilon}{2}F_{\mu\nu}X^{\mu\nu}+\frac{1}{2}m_{X}^{2}X_{\mu}X^{\mu}$ (1)
where $F_{\mu\nu}$ and $X_{\mu\nu}$ are the field strength tensors for the SM photon and the DP, $m_{X}$ is the DP mass, and $\varepsilon$ is the kinetic mixing parameter. There are several interesting DP production mechanisms that may address the correct abundance of DM. One of the simplest ways to produce DP is the misalignment mechanism Nelson and Scholtz (2011), another one is based on the tachyonic instability arising when the DP couples to a misalignment axion Co et al. (2019), and a third one is from the decay of cosmic strings Long and Wang (2019). DP on the Earth, may come from the burning of the Sun with the DP flux proportional to the photon flux rescaled by a factor $\varepsilon^{2}$, or from the relic density of DM, which is about $0.4~{}{\rm GeV/cm^{3}}$. The absorption of DP by a superconductor deposits energies in the term of phonon, which may subsequently break the Cooper pair. This effect will be investigated in the next section.
@@
where $m_{f}$ is the fermion mass. Given these interactions, an ALP can be absorbed by the electron, proton or neutron when it pass through a target, producing phonon signal. Free parameters in this process includes: $m_{a}$, $g_{a}$ and the ALP flux. The ALP flux may come from the cosmic relic abundance controlled by $\rho_{\rm DM}$ and its velocity distribution. Alternatively it can be produced in stars Carenza et al. (2024), especially in the sun and other nearby stars. These flux will be applied to the following analysis.
@@ III Detectability at transmon qubits
World-wide efforts toward “quantum advantage” Arute et al. (2019); Zhong et al. (2020); Wu et al. (2021) has intensified the competition in quantum information devices, such as circuit quantum electrodynamics devices Blais et al. (2021). Among these devices, the transmon qubit is one of the most widely used in various quantum computing architectures. A transmon qubit Ristè et al. (2013), consisting of a large shunting capacitor connected in parallel with the Josephson junction Makhlin et al. (2001), is a modified version of the charge qubit, aka a “Cooper-pair box”. The trasmon qubit has a long coherence time that is essential for quantum computation. Its resistance to charge noise makes it an ideal device for dark matter detections Chen et al. (2023, 2024); Das et al. (2024a, b); Moretti et al. (2024).
In this section we follow the strategy of Refs. Das et al. (2024a, b) to investigate the indirect detection sensitivity by using a transmon qubit to absorb a low mass DM particle. Ref. Ristè et al. (2013) investigated the decoherence of a transmon qubit, using a single-junction superconducting qubit made by aluminum. A quasi-particle density is measured as $0.04\pm 0.01~{}\mu m^{-3}$. It is well-known that low temperature allows electrons to form Cooper pairs, bound via long range interactions with phonons. When DM scatters with the superconducting aluminum, its energy is partially deposited in the form of phonons, which can break Cooper pairs if phonons have energy exceeding the binding energy of Cooper pairs and lead to an excess of quasiparticles. The steady-state quasiparticle density $n_{qp}$ can be estimated using the mean field results Bespalov et al. (2016),
@@
where $\Gamma_{R}$, $\Gamma_{T}$ and $\Gamma_{G}$ ar the recombination, trapping and generation rates, respectively. $\Gamma_{R}\sim\bar{\Gamma}n_{QP}^{2}$ and $\Gamma_{T}=\bar{\Gamma}_{T}n_{\rm QP}$, while $\Gamma_{G}$ can be related to the scattering of DM-off the superconductor. $\bar{\Gamma}=40~{}\mu m^{3}/s$ for aluminium Bespalov et al. (2016) and $\bar{\Gamma}_{T}$ is set to be zero for simplicity.
For quasiparticles in equilibrium, one has
@@
which can be used to determine the DM-superconductor scattering rate or absorption rate given the value of $n_{\rm QP}$. On the other hand, the quasiparticle generation rate can be written as Das et al. (2024a, b)111 Here we have included an extra factor $\rho_{T}$ in the formula, considering that $R$ is DM interaction rate per unit time per unit target mass.
@@
where $\varepsilon=0.6$ Hochberg et al. (2023); Kaplan et al. (1976) being the quasi-particle generation efficiency, and $\Delta=340~{}{\mu eV}$ denotes the superconducting energy gap of aluminum, $R$ is DM scattering or absorption rate per unit time per unit target mass. Combing Eqs. (6) and (5), one may derive a constraint on the DM coupling strength for a given DM scattering rate or absorption rate. In the following, we will consider low-mass DM candidates including an ALP, a scalar DM and the dark photon, and derive their respective detection limits.
@@ III.1 Dark photon
Given the discussion above, we can place constraints on different types of bosonic DM by measuring quasiparticles with a transmon qubit. First, let us calculate the energy deposition by light bosonic dark matter. Recent studies  Hardy and Lasenby (2017); Mitridate et al. (2021); Chen et al. (2022); Krnjaic and Trickle (2023); Mitridate et al. (2024) have shown that the DM absorption rate can be written in terms of in-medium self-energies. by the optical theorem, the absorption rate of the $\lambda$th polarization of the DM, $\chi$, can be written as
$\Gamma^{\lambda}=-\frac{1}{m_{\chi}}{\rm Im}[\Pi_{\chi\chi}^{\lambda}]$ (7)
where $\Pi_{\chi\chi}^{\lambda}$ is the self-energy between two $\chi$ particles of the $\lambda$th polarization. $\Pi_{\chi\chi}^{\lambda}$ receive contributions from both electronic excitations and phonon excitations. Assuming the electronic band gap is much larger than the energy of phonon excitations and taking use of real part of the electron self-energy given in Refs Mitridate et al. (2021); Chen et al. (2022); Krnjaic and Trickle (2023), one only needs to consider phonon contribution to self energies, which has been studied in the Ref. Mitridate et al. (2024).
As a result, the total absorption rate per unit target mass per unit time can be written as
@@
where $\rho_{\rm DP}$ is the local DP energy density. $\rho_{T}$ is the target energy density that depends on the material. For aluminum at $T_{C}=1.2$ K, $\rho_{T}$ is typically $2.7k/cm^{3}$. $\kappa$ is the DP-photon kinetic mixing parameter, and the last term is the energy loss function (ELF), which is evaluated using the public code DarkELF Knapen et al. (2022).
@@
where $\nu_{0}=1.2\times 10^{4}~{}\mu m^{-3}\mu eV^{-1}$ being the Cooper pair density at the Fermi level in the aluminum. It means that the curve in the plot only gives the exclusion limit on the parameter space: regime above the curve is excluded. Alternatively, if one can precisely evaluate the $n_{qp}^{th}$, one can take $n_{qp}^{\rm DM}$ as a distinctive signal of new physics beyond the SM, and a possible signal of DM.
Figure 2: Left-panel: Transmon qubit constraint on the light scalar DM, where the light-blue regime is excluded by the fifth force measurement Adelberger et al. (2003) and horizontal line is the limit of the White Dwarfs observations Bottaro et al. (2023). Right-panel: Transmon qubit constraint on the axion-like particle in the $m_{a}-g_{aee}$ plane (solid), based on a sensitivity estimate with $n_{\rm qp}\leq 0.04~{}{\rm\mu m}^{-3}$. The current limit from measurements on the electron dipole moment $g_{e}-2$ is shown for comparison (red shaded).

### v3

# Probing Sub-eV Dark Photon, Scalar and Axion-like Particle Dark Matters with Transmon Qubits

@@ Abstract
In this paper, we investigate constraints of the transmon qubit, an improved version of the charge qubit, on bosonic light dark matters. Phonon excitations induced by the scattering or absorption of dark matter on a superconductor may destroy the Cooper pair, leading to the production of quasiparticles made by the electron. By measuring the production rate of the quasiparticle density, one may read out the coupling between dark matter and ordinary matter, assuming that these quasiparticles are solely induced by dark matter interactions. For the first time, we show constraints on the parameter space of the dark photon, light scalar dark matter, and axion-like particles from the measurement of quasiparticles in transmon qubit experiments. This study offers insights for the development of quantum qubit experiments aimed at the direct detection of dark matter in underground laboratories.
@@ I Introduction
In recent years, direct detection technologies for DM Billard et al. (2022); Schumann (2019); Marrodán Undagoitia and Rauch (2016); Fitzpatrick et al. (2013); Essig et al. (2012) have developed rapidly. The exclusion limits provided by experiments utilizing nuclear recoil technology have been continuously strengthened, and numerous experiments employing electronic excitation have been conducted to probe DM, enabling the detection of lighter DM candidates. However, these experiments are mostly limited to non-relativistic DM candidates with a mass above the MeV scale. Furthermore, when considering the DM absorption process, the detection threshold can be lowered to the electroVolt (eV) level. In fact, there are also sub-eV DM models with strong theoretical motivations, such as the axion and the dark photon. These particles can be produced in the early universe through mechanisms like the misalignment mechanism and parametric resonance, etc. Direct Detection of ultralight DM Hui et al. (2017); Dror and Verner (2025); Fukuda et al. (2019); Sun et al. (2025) is an important unresolved issue, and various proposals have been put forward to explore sub-eV DM. Electronic excitations in targets with small excitation gaps can be applied to the detection of light DM Essig et al. (2017). The expected target materials include Dirac materials Hochberg et al. (2018), superconductors Hochberg et al. (2016), and narrow-gap semiconductors Hochberg et al. (2017). The emergence of collective excitations like phonon Kahn and Lin (2022); Augier et al. (2023); Campbell-Deem et al. (2022); Lyon et al. (2024); Coskuner et al. (2022); Cox et al. (2019); Chao et al. (2024), magnon Trickle et al. (2020); Mitridate et al. (2020); Chigusa et al. (2020, 2024); Barbieri et al. (1989); Flower et al. (2019) and plasmon Liang et al. (2025); Liu et al. (2023); Kozaczuk and Lin (2020), resulting from DM-matter scattering or DM absorption within condensed matter material, represents a promising avenue for DM direct detection. This is particularly relevant given that these excitations possess energies within the ${\cal O}(1\sim 100)$ meV range, aligning well with the energy scales characteristic of light DM. In brief, researchers are actively exploring and employing a variety of methods to detect DM, yet no definitive detection has been achieved. Utilizing new materials or technologies to detect light DM presents an increasingly recognized challenge in the fields of high-energy physics and precision measurement physics. The path to the direct detection of DM remains a lengthy and challenging journey.
@@
In this study, we delve into the potential constraints imposed on light DM by superconducting qubits Dixit et al. (2021). These quantum qubits, constructed from superconducting materials and maintained at ultra-low temperatures, have garnered significant interest among quantum computing scientists for their enhanced coherence and stability. Recently, transmon qubits Ristè et al. (2013) as DM detectors have been proposed. Refs. Chen et al. (2023, 2024) have showed how to detect hidden photon and the QCD axion using the direct excitation of transmon qubits. Refs. Das et al. (2024a, b) have shown constraint on the DM-nucleon scattering cross sections using the latest transmon qubit measurements Dixit et al. (2021). In addition, a transmon qubit prototype for DM detection experiment has been designed in Ref. Moretti et al. (2024). Compared with conventional direct detection experiments, the transmon qubit facility have several merits on the direct detection of light DM: (1) Transmon qubits maintain long quantum coherence up to microseconds at cryogenic temperatures, enabling precise detection of faint energy signals from DM interactions; (2) Operated in dilution refrigerators, transmon qubits benefit from minimal thermal and electromagnetic noise, suppressing interference that could obscure DM signals, which is critical for detecting DM interactions with extremely low energy threshold; (3) The energy level of transmon qubits can be precisely adjusted using external microwaves or magnetic fields, adapting to different DM models with specific couplings. In addition, arrays of transmon qubits on a single chip can simultaneously detect DM signals across multiple frequency bands, enhancing detection efficiency and statistical significance. All these advantages make transmon qubit experiments a potential DM direct detection approach worthy of continuous investigation, especially for probing light DM. In this paper, We focus on the case of the absorbable DM. Light bosonic DM can be absorbed by the superconducting material into phonons via DM-proton or DM-neutron interactions and the absorption rate can be calculated utilizing the effective field theory developed in Ref. Mitridate et al. (2024). These phonon can break Cooper pairs whenever their energy is above meV, leading to to the production of quasiparticles. With the help of transmon qubit measurements Dixit et al. (2021), we put constraints on the DM couplings with results shown in Figs. 1 and 2, where constrains on the dark photon, light scalar DM and axion-like particle DM are presented. These results are the first solid constraints on light bosonic DM from quantum qubits experiments, and will insights for the development of quantum qubit experiments aimed at the direct detection of dark matter in underground laboratories.
@@ II Light Bosonic DMs
${\cal L}_{\rm DP}\supset-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}F^{\prime}_{\mu\nu}F^{\prime\mu\nu}+\frac{\kappa}{2}F_{\mu\nu}F^{\prime\mu\nu}+\frac{1}{2}m_{X}^{2}X_{\mu}X^{\mu}$ (1)
where $X_{\mu}$ is the DP field, $F_{\mu\nu}$ and $F^{\prime}_{\mu\nu}(=\partial_{\mu}X_{\nu}-\partial_{\nu}X_{\mu})$ are the field strength tensors for the SM photon and the DP, $m_{X}$ is the DP mass, and $\kappa$ is the kinetic mixing parameter. There are several interesting DP production mechanisms that may address the correct abundance of DM. One of the simplest ways to produce DP is the misalignment mechanism Nelson and Scholtz (2011), another one is based on the tachyonic instability arising when the DP couples to a misalignment axion Co et al. (2019), and a third one is from the decay of cosmic strings Long and Wang (2019). DP on the Earth, may come from the burning of the Sun with the DP flux proportional to the photon flux rescaled by a factor $\kappa^{2}$, or from the relic density of DM, which is about $0.4~{\rm GeV/cm^{3}}$. The absorption of DP by a superconductor deposits energies in the term of phonon, which may subsequently break the Cooper pair. This effect will be investigated in the next section.
@@
where $m_{f}$ is the fermion mass. Given these interactions, an ALP can be absorbed by the electron, proton or neutron when it pass through a target, producing phonon signal. Free parameters in this process includes: $m_{a}$, $g_{aff}$ and the ALP flux. The ALP flux may come from the cosmic relic abundance controlled by $\rho_{\rm DM}$ and its velocity distribution. Alternatively it can be produced in stars Carenza et al. (2024), especially in the sun and other nearby stars. In this study we will not consider DM flux emitted from the sun, whose impact to the quasi-particle will be studied in a future work. On the other hand we take the dark matter flux as $\rho_{\rm DM}v$ with $v$ the velocity of DM and $\rho_{\rm DM}(\sim 0.4~GeV/cm^{3})$ the local DM energy density.
@@ III Detectability at transmon qubits
In recent years, the global pursuit of quantum advantage Arute et al. (2019); Zhong et al. (2020); Wu et al. (2021) has escalated the competition in the development of quantum calculation devices. This race encompasses a diverse range of technologies, with circuit quantum electrodynamics (cQED) devices [56] at the forefront of research and innovation. Among these devices, the transmon qubit has emerged as one of the most pivotal components across various quantum computing architectures, owing to its remarkable performance and scalability. A transmon qubit Ristè et al. (2013) represents an advanced iteration of the charge qubit, also known as a Cooper-pair box. Its design features a Josephson junction—a fundamental element consisting of two superconducting layers separated by a thin insulating barrier—connected in parallel with a large shunting capacitor Makhlin et al. (2001). This configuration addresses the inherent limitations of traditional charge qubits, particularly their susceptibility to charge noise. By introducing the shunting capacitor, the transmon qubit effectively suppresses fluctuations in electrical charge, thereby stabilizing its quantum state. One of the transmon qubit’s improtant attributes is its extended coherence time, a parameter that quantifies how long the qubit can maintain its quantum state without decohering or losing information. This characteristic makes transmon qubits indispensable for building practical quantum computers capable of tackling complex problems beyond the reach of classical machines. Beyond quantum computing, the transmon qubit’s resilience to charge noise makes it as an ideal candidate for DM direct detection experiment Chen et al. (2023, 2024); Das et al. (2024a, b); Moretti et al. (2024). Transmon qubits can serve as highly sensitive detectors, leveraging their ability to detect minute energy exchanges or quantum state perturbations that may result from interactions with DM particles. Their low-noise operation and precise quantum state control enable the identification of faint signals that would otherwise be obscured by background noise, thus opening up new avenues for exploring the nature of the universe’s elusive DM component.
In this section we follow the strategy of Refs. Das et al. (2024a, b) to investigate the indirect detection signal of DM by measuring the quasi-particle density induced by phonons which come from the scattering or absorption a low mass DM particle on the facility. Ref. Ristè et al. (2013) has investigated the decoherence of a transmon qubit, using a single-junction superconducting qubit made by aluminum, and the quasi-particle density is measured as $n_{qp}\approx 0.04\pm 0.01~\mu m^{-3}$. It is well-known that low temperature allows electrons to form Cooper pairs, bound via long range interactions with phonons. When DM scatters with the superconducting aluminum, its energy is partially deposited in the form of phonons, which can break Cooper pairs if phonons have energy exceeding the binding energy of Cooper pairs and lead to an excess of quasiparticles. The steady-state quasiparticle density $n_{qp}$ can be estimated using the mean field results Bespalov et al. (2016),
@@
where $n_{qp}$ is the quasi-particle density, $\Gamma_{R}$, $\Gamma_{T}$ and $\Gamma_{G}$ are the recombination rate, trapping rate and generation rate, respectively. The recombination and trapping rates can be written as $\Gamma_{R}\sim\bar{\Gamma}n_{QP}^{2}$ and $\Gamma_{T}=\bar{\Gamma}_{T}n_{\rm QP}$, while $\Gamma_{G}$ is related to the interaction of DM with the superconductor material, which will be discussed in the following. In our analysis, we take $\bar{\Gamma}=40~\mu m^{3}/s$ for aluminium Bespalov et al. (2016) and take $\bar{\Gamma}_{T}=0$ Das et al. (2024a, b), for simplicity.
For quasiparticles in equilibrium, one has $dn_{pq}/dt=0$, and thus
@@
which can be used to determine the DM-superconductor scattering rate or absorption rate given the value of $n_{\rm QP}$. On the other hand, the quasiparticle generation rate can be written as Das et al. (2024a, b) [^1] Here we have included an extra factor $\rho_{T}$ in the formula, considering that $R$ is DM interaction rate per unit time per unit target mass.
$\Gamma_{G}=\frac{\zeta}{2\Delta}\int d\omega\omega\frac{dR}{d\omega}\rho_{T}$ (6)
where $\zeta=0.6$ Hochberg et al. (2023); Kaplan et al. (1976) being the quasi-particle generation efficiency, and $\Delta=340~{\mu eV}$ denoting the superconducting energy gap of aluminum, $R$ is DM scattering or absorption rate per unit time per unit target mass, $\omega$ is the frequency of the phonon and $\rho_{T}$ is the density of the target. Combing Eqs. (6) and (5), one may derive a constraint on the DM coupling strength for a given DM scattering rate or absorption rate. In the following, we will consider constraints of transmon qubit measurements on low-mass DM candidates including an ALP, a scalar DM and a dark photon, by deriving limits on their couplings.
@@
$\Gamma^{\lambda}=-\frac{1}{m_{\chi}}{\rm Im}[\Pi_{\chi\chi}^{\lambda}]$ (7)
[…]
@@ III.1 Dark photon
As a result, the total absorption rate of the dark photon dark matter per unit target mass per unit time can be written as
@@
where $\rho_{\rm DP}$ is the local dark photon energy density. $\rho_{T}$ is the target energy density that depends on the material. For aluminum, $\rho_{T}$ is typically $2.7g/cm^{3}$. $\kappa$ is the DP-photon kinetic mixing parameter given in the Eq. (1), and the last term is the energy loss function (ELF), which is evaluated using the public code DarkELF Knapen et al. (2022), $\varepsilon(\omega)$ is the frequency dependent longitudinal dielectric function.
@@
where $\nu_{0}=1.2\times 10^{4}~\mu m^{-3}\mu eV^{-1}$ being the Cooper pair density at the Fermi level in the aluminum, $k_{B}$ is the Boltzmann constant. It means that the curve in the plot only gives the exclusion limit on the parameter space: regime above the curve is excluded. It should be mentioned that the thermal quasiparticle density at a temperature $T$ is indeed a key parameter setting the baseline for any detection experiment. Our assumptions for the thermal background are conservative and align with these established experimental results. We fully agree that any future dedicated dark matter detector would require in-situ characterization of its specific quasiparticle background, the constraints we derive are based on experimentally validated physics. The projected sensitivity for a future optimized device would, of course, be contingent upon achieving thermal quasiparticle densities at or below the levels already demonstrated in state-of-the-art qubit experiments.
Figure 2: Left-panel: Transmon qubit constraint on the light scalar DM, where the light-blue regime is excluded by the fifth force measurement Adelberger et al. (2003) and horizontal line is the limit of the White Dwarfs observations Bottaro et al. (2023). Right-panel: Transmon qubit constraint on the axion-like particle in the $m_{a}-g_{aee}$ plane (solid), based on a sensitivity estimate with $n_{\rm qp}\leq 0.04~{\rm\mu m}^{-3}$. The limit from measurements on the electron dipole moment $g_{e}-2$ is shown for comparison (red shaded). Other constraints are not included considering that $g_{e}-2$ results encompass all constraints of transmon qubit.